### 14 Apr 2026 v2

# Optical pumping of alkali-metal vapor with hyperfine-resolved buffer gas pressure

Abstract
Optical pumping is fundamental to high-precision measurement using thermal alkali-metal atoms in vapor cells. In applications such as spin-exchange-relaxation-free magnetometers, buffer gases (e.g., $\text{N}_{2}$ or He) are commonly employed to quench fluorescence and mitigate wall relaxation. In the high-pressure limit (e.g., the $\text{N}_{2}$ pressure $p_{\rm{N}_{2}}>1$ atm), where collisional broadening exceeds the hyperfine splitting of the alkali-metal atoms, optical pumping theory provides a clear description of the angular momentum exchange between photons and atomic spins. However, in many magnetic sensing scenarios, this high-pressure condition is not strictly satisfied, rendering the high-pressure approximation inaccurate. Consequently, a precise quantitative understanding of optical pumping under realistic pressures is critical for determining optimal buffer gas parameters, selecting operating points (e.g., pump frequency and intensity), and enhancing system reliability and stability. To address this, we develop a theory of optical pumping in the quasi-high-pressure regime, where collisional broadening is comparable to the ground-state hyperfine splitting. We demonstrate that optical absorption, spin polarization, and magnetic resonance linewidth in this regime differ significantly from those predicted by the high-pressure limit and offer favorable operating conditions. Our study extends conventional modeling and offers critical guidance for atomic magnetometry operating under realistic buffer gas pressures.
I Introduction
Optical pumping of alkali-metal atoms serves as the foundation for a wide array of quantum sensing technologies, ranging from atomic clocks [14, 23] to magnetometers [4, 3, 7] and inertial sensors [28, 26, 8]. In particular, atomic magnetometers exploit the spin dynamics of alkali-metal atoms to achieve ultrasensitive field detection. In this process, a circularly polarized light beam creates atomic spin polarization, which then precesses at the Larmor frequency in an external magnetic field, providing a precise measure of the field strength.
Vapor cells used in these applications are generally categorized as either evacuated or buffer-gas-filled [12]. While evacuated cells rely on anti-relaxation coatings to prevent wall-collision relaxation [21, 20], buffer-gas-filled cells utilize inert gases to modify atomic diffusion and quench fluorescence [25, 17]. These features allow operation at higher temperatures and atomic densities, enabling the ultra-high sensitivity characteristic of modern magnetometers [18, 15]. This work focuses specifically on the optical pumping dynamics within such buffer-gas-filled cells.
Depending on the buffer gas pressure, optical pumping operates in distinct regimes, yet a comprehensive theoretical framework covering the intermediate pressures is currently lacking. In the low-pressure (LP) regime with typical buffer gas pressure $\lesssim 10~\mathrm{Torr}$, strong and narrow optical resonances of well resolved hyperfine mutiplets enables refined optical manipulation, such as in atomic clocks based on coherent population trapping (CPT) [24]. In the high-pressure (HP) limit (e.g., buffer gas pressure $1~\mathrm{atm}$), strong collisional quenching and pressure broadening dominate [9, 27, 22]. This treatment renders hyperfine mutiplets unresolved and forms the basis of the spin-temperature distribution (STD) widely used in, e.g., spin-exchange relaxation free (SERF) magnetometers [11, 19]. However, many optical pumping systems including atomic magnetometers operate in the quasi-high-pressure (QHP) regime (typically $\sim 10^{2}~\mathrm{Torr}$) [15, 5]. We focus on this intermediate regime, in which the ground-state (GS) hyperfine structure remains spectrally resolvable while the excited-state (ES) structure is unresolved. This leads to spin dynamics that are neither fully captured by the LP models nor the simplified HP approximations. Developing a precise quantitative understanding of the QHP regime is therefore critical. It not only enables the determination of optimal operating points to maximize sensitivity, but also facilitates the assessment of long-term system stability against parameter drifts.
The spectral resolvability of the GS hyperfine structure fundamentally governs the optical pumping process in the QHP regime. Unlike the HP limit, the QHP regime requires a rigorous treatment of the resolvable multiplets and their distinct contributions to absorption. Starting from the master equation in Liouville space, we incorporate the electric-dipole interaction and relevant collisional mechanisms to derive the mathematical form of optical pumping in the QHP regime. Crucially, we establish a compact relation linking QHP superoperators to their HP counterparts, providing a unified theoretical framework that facilitates both physical analysis and efficient numerical evaluation.
A distinct feature of the QHP regime is the hyperfine-resolved dependence of absorption cross-section on pumping-light polarization and intensity, which markedly differs from the HP limit. Within our framework, we derive analytical expressions for absorption under both linearly and circularly polarized pumping. We elucidate how the redistribution of population among and within hyperfine multiplets reshapes the detuning response of the absorption behavior.
Consequently, driven by the combined redistribution dynamics in the QHP regime, the steady-state atomic population generally deviates from the standard STD. We demonstrate that while the STD can approximate the distribution shape under rapid spin-exchange collisions, the effective temperature parameter—namely the spin polarization—requires QHP-specific corrections. Evaluating these dynamics, we demonstrate that resonant pumping at specific hyperfine levels can yield a larger polarization.
Beyond fundamental spin dynamics, we analyze the magnetic resonance linewidth to identify optimal operating conditions for atomic magnetometers. We find that resonant pumping at the lower hyperfine-split levels not only enhances spin polarization but also yields a favorable magnetic-resonance linewidth. This suggests a strategy for optimizing sensitivity through simultaneous improvements in resonance amplitude and linewidth.
The remainder of this paper is organized as follows. Section II defines the QHP regimes based on spectral characteristics. Section III presents the QHP-specific theoretical framework based on the Liouville space master equation. Section IV discusses the unique features of absorption, spin polarization, and magnetic resonance linewidth in the QHP regime. Finally, Section V summarizes the main findings.
II Regimes of buffer gas pressure
Figure 1: Schematic illustration of optical pumping on the alkali-metal D1 line in the HP limit and QHP regime. (a) and (b) show the GS and ES hyperfine multiplets, with different collisional broadening $\Gamma_{\mathrm{brd}}$ illustrated by blue level edges. Each of $F$ and $\bar{F}$ corresponds to either $a=I+1/2$ or b=I1/2b=I-1/2 for nuclear spin II. Laser detuning from the D1 line reference frequency is denoted by $\Delta_{0}$, while $\Delta_{F}$ for $F=a$ and $b$ are the detunings corrected from $\Delta_{0}$ by the GS hyperfine splitting. (c) and (d) present the corresponding absorption cross section $\sigma_{0}$ for unpolarized 87Rb, with $\Gamma_{\mathrm{brd}}=3,\,20~\mathrm{GHz}$, respectively. The hyperfine splittings are $\omega^{\mathrm{\{g\}}}_{\mathrm{hf}}=6.8~\mathrm{GHz}$ and $\omega^{\mathrm{\{g\}}}_{\mathrm{hf}}=0.81~\mathrm{GHz}$. Collisional shifts are not included for simplicity. Absorption peaks corresponding to different $F\rightarrow\bar{F}$ transitions are indicated. Doppler broadening $\Gamma_{\mathrm{dop}}=0.57~\mathrm{GHz}$ is included using the Voigt profile, which is negligible in the two regimes.
Optical pumping in alkali-metal vapor relies on the coupling between GS and ES of the atomic D line transition, which drives atoms toward a targeted state via optical excitation and collisional relaxation. The $\mathrm{D}_{1}$ line typically serves as the model system, as shown in Fig. 1, where $F$ and $\bar{F}$ denote the angular momentum of the GS and ES hyperfine multiplets, respectively. While light absorption initiates the pumping process, the spectral characteristics are governed by line broadening mechanisms. In buffer-gas vapor cells, which are the focus of this work, collisional broadening $\Gamma_{\mathrm{brd}}$ dominates, serving as the primary criterion for distinguishing pressure regimes.
To characterize these regimes, we analyze the absorption cross-section for unpolarized atoms, $\sigma_{0}$, in the limit of vanishing photon flux $\Phi$ of the incident laser beam (i.e., $\Phi\rightarrow 0$). In the LP regime, $\Gamma_{\mathrm{brd}}$ is insufficient to bridge the ES or GS hyperfine splitting, and dynamics are further complicated by radiation trapping [17] and atomic diffusion [13]. Instead of this regime, we focus on the distinction between the HP limit and the QHP regime in the following. In the HP limit, $\Gamma_{\mathrm{brd}}$ is sufficiently large to merge all GS and ES hyperfine splittings into a single resonance peak, as depicted in Figs. 1(a) and 1(c). Here, the system simplifies significantly, and underlies the STD [11, 19].
The QHP regime represents an intermediate condition where $\Gamma_{\mathrm{brd}}$ is comparable to the GS hyperfine splitting but significantly larger than that of the ES. Figures 1(b) and 1(d) illustrate that the absorption profile simplifies to two resolved components associated with the GS multiplets $F=a$ and $b$. Unlike the HP limit, where atoms in different multiplets absorb photons at an almost common rate, in the QHP regime, the pumping process becomes more sensitive to the pump laser frequency. Furthermore, stronger pump light significantly alters the population distribution between the =a=a and $b$ levels, which in turn reshapes the absorption behavior and spin polarization—factors critical for subsequent applications.
While the linear absorption profile in the $\Phi\rightarrow 0$ limit serves as a convenient classification metric, the rich physical phenomena emerging at higher optical intensities are the primary focus of this work. We emphasize that the behaviors in the QHP regime cannot be fully described by the simple optical pumping model used in the HP limit. Therefore, a key contribution of this paper is the development of a unified theoretical framework capable of treating the continuous transition from the HP limit to the QHP regime. The following sections will present this unified approach and analyze the spin evolution in practical scenarios, such as atomic magnetometers, where operation often spans these intermediate pressure conditions.
III THEORETICAL TREATMENT
III.1 Master equation in Liouville space
We study the spin dynamics of optically pumped alkali-metal atoms in vapor cells, where both coherent and incoherent processes contribute to the evolution of the atomic ensemble. For the coherent part, the atomic spin Hamiltonian is [12]
$H_{0}=$ $\frac{\omega^{\{\mathrm{eg}\}}}{2}\left(1^{\{\mathrm{e}\}}-1^{\{\mathrm{g}\}}\right)+H^{\{\mathrm{g}\}}+H^{\{\mathrm{e}\}},$ (1)
$H^{\{\mathrm{g}\}}=$ $\sum_{F,m_{F}}\omega^{\{\mathrm{g}\}}_{F,m_{F}}|F,m_{F}\rangle\langle F,m_{F}|,$ (2)
$H^{\{\mathrm{e}\}}=$ $\sum_{\bar{F},\bar{m}_{F}}\omega^{\{\mathrm{e}\}}_{\bar{F},\bar{m}_{F}}|\bar{F},\bar{m}_{F}\rangle\langle\bar{F},\bar{m}_{F}|,$ (3)
where $1^{\{\mathrm{g}\}}$ and $1^{\{\mathrm{e}\}}$ are identity operators of the GS and ES subspaces, respectively, and $\omega^{\{\mathrm{eg}\}}$ denotes the reference frequency of the optical transition. In Eqs. (2) & (3), the Hamiltonians $H^{\{\mathrm{g}\}}$ and $H^{\{\mathrm{e}\}}$ are expressed in their eigen-representations with eigen-frequencies $\omega^{\{\mathrm{g}\}}_{F,m_{F}}$ and $\omega^{\{\mathrm{e}\}}_{\bar{F},\bar{m}_{F}}$, together with the eigen-states $|F,m_{F}\rangle$ and $|\bar{F},\bar{m}_{F}\rangle$, respectively. In this paper, we assume the magnetic field is so weak that the total angular momentum $F$ and $\bar{F}$ are approximately good quantum numbers, with corresponding magnetic quantum numbers $m_{F}$ and $\bar{m}_{F}$.
Having specified the atomic Hamiltonian $H_{0}$, we introduce the light-atom interaction Hamiltonian $H_{\mathrm{int}}$, which takes an electric-dipole form [12]
$H_{\mathrm{int}}=V+V^{\dagger}=-\Omega_{\mathrm{I}}\bm{d}^{\dagger}\cdot\hat{\bm{e}}-\Omega^{*}_{\mathrm{I}}\bm{d}\cdot\hat{\bm{e}}^{*},$ (4)
with $V$ describing photon absorption and atom excitation with optical Rabi frequency $\Omega_{\mathrm{I}}$, and $V^{\dagger}$ representing the reverse process. Here, $\bm{d}$ denotes the dimensionless dipole operators, and $\hat{\bm{e}}$ is the unit vector of the positive-frequency classical electric field. For a monochromatic laser beam with frequency $\omega$, the laser detuning with respect to the atomic transition $|F,m_{F}\rangle\leftrightarrow|\bar{F},\bar{m}_{F}\rangle$ is
[…]
where $\Delta_{0}=\omega-\omega^{\{\mathrm{eg}\}}$ is the hyperfine-splitting-independent detuning of the laser frequency $\omega$ to the reference frequency of the optical transition $\omega^{\{\mathrm{eg}\}}$.
In addition to the coherent mechanisms governed by the Hamiltonians $H_{0}$ and $H_{\mathrm{int}}$, the atoms experience incoherent dynamics, including level broadening and spin relaxation due to atomic collisions, spontaneous emission of photons, and non-radiative decay (quenching) processes. Due to these incoherent processes, the optical ES are hardly occupied. In this case, the ES degree of freedom and the optical coherences are adiabatically eliminated [12]. Therefore, the evolution of atomic spin is reduced to the master equation in GS Liouville space [1]
$\frac{d}{dt}\left|{\rho}\right)=$ $-\left(iH^{\{\mathrm{g}\}\copyright}+\mathcal{G}^{\{\mathrm{g}\}}_{\mathrm{rel}}+R_{\mathrm{op}}\mathcal{A}_{\mathrm{op}}\right)\left|\rho\right),$ (6)
where $\left|\rho\right)$ is the vectorized GS density matrix, and $X^{\copyright}=X^{\flat}-X^{\dagger\sharp}$ is the commutator superoperator with $X^{\flat}$ and $X^{\sharp}$ the left- and right-translation superoperators [9], corresponding to the operator $X$ in Hilbert space. The GS spin relaxation term $\mathcal{G}^{\{\mathrm{g}\}}_{\mathrm{rel}}$ is
$\mathcal{G}^{\{\mathrm{g}\}}_{\mathrm{rel}}$= $\Gamma_{\mathrm{SD}}\mathcal{A}_{\mathrm{SD}}+\Gamma_{\mathrm{ex}}\left(\mathcal{A}_{\mathrm{SD}}-(\mathbf{S}|{\rho})\cdot\bm{\mathcal{A}}_{\mathrm{SE}}\right),$ (7)
with the spin-destruction rate $\Gamma_{\mathrm{SD}}$, the spin-exchange rate $\Gamma_{\mathrm{ex}}$, and the expectation value of the electron spin vector $(\mathbf{S}|\rho)=\mathrm{Tr}\left[\mathbf{S}{\rho}\right]$. The spin-destruction and spin-exchange superoperators $\mathcal{A}_{\mathrm{SD}}$ and $\mathcal{A}_{\mathrm{SE}}$ are
$\mathcal{A}_{\mathrm{SD}}=$ $\frac{3}{4}-\bm{\mathbf{S}}^{\flat}\cdot\bm{\mathbf{S}}^{\sharp},$ (8)
$\bm{\mathcal{A}}_{\mathrm{SE}}=$ $\bm{\mathbf{S}}^{\flat}+\bm{\mathbf{S}}^{\sharp}-2i\bm{\mathbf{S}}^{\flat}\times\bm{\mathbf{S}}^{\sharp},$ (9)
where $\mathbf{S}$ is the electron spin operator of the alkali-metal atom.
The optical pumping process is described by the third term of the right-hand-side of Eq. (6), where $R_{\mathrm{op}}$ is the optical pumping rate and $\mathcal{A}_{\mathrm{op}}$ is the corresponding superoperator. The optical pumping process consists of two parts [12, 9]
$\mathcal{A}_{\mathrm{op}}=\mathcal{A}_{\mathrm{dp}}-\mathcal{A}_{\mathrm{rp}},$ (10)
where $\mathcal{A}_{\mathrm{dp}}$ and $\mathcal{A}_{\mathrm{rp}}$ represent the depopulation and repopulation processes, respectively. The depopulation term $\mathcal{A}_{\mathrm{dp}}$ is associated with light absorption and removes population from the GS,
$R_{\mathrm{op}}\mathcal{A}_{\mathrm{dp}}=$ $i[V^{\dagger}W]^{\copyright}\equiv i[\delta H^{\{\mathrm{g}\}}]^{\copyright},$ (11)
where
$W=$ $V./(\Delta^{\{\mathrm{eg}\}}+i\Gamma_{\mathrm{brd}}).$ (12)
In Eq. (12), $\Delta^{\{\mathrm{eg}\}}=\left[\Delta_{\bar{F}\bar{m}_{F},Fm_{F}}\right]$ is the detuning matrix with elements defined in (5), and the dot–slash product is defined by $\langle\nu|A./B|\mu\rangle=\langle\nu|A|\mu\rangle/\langle\nu|B|\mu\rangle$ for matrices $A$ and $B$ [9]. The repopulation term $\mathcal{A}_{\mathrm{rp}}$, describing the optical excitation, spin relaxation, and quenching of the atomic ES and restoring the GS, is shown to be the form [1]
$\mathcal{A}_{\mathrm{rp}}=$ $\left(\Gamma_{\mathrm{s}}\mathcal{A}^{\{\mathrm{ge}\}}_{\mathrm{s}}+\Gamma_{\mathrm{q}}\mathcal{A}_{\mathrm{q}}^{\{\mathrm{ge}\}}\right)$
$\times\left[{iH^{\{\mathrm{e}\}\copyright}}+\Gamma_{\mathrm{s}}+\Gamma_{\mathrm{q}}+\Gamma_{\mathrm{JD}}{A}_{\mathrm{JD}}\right]^{-1}\mathcal{A}_{\mathrm{p}}^{\{\mathrm{eg}\}},$ (13)
where $\mathcal{A}_{\mathrm{s}}^{\{\mathrm{ge}\}}$ and $\mathcal{A}_{\mathrm{q}}^{\{\mathrm{ge}\}}$ are the superoperators of the spontaneous emission and quenching processes with corresponding rates $\Gamma_{\mathrm{s}}$ and $\Gamma_{\mathrm{q}}$. The superaoperator $\mathcal{A}_{\mathrm{JD}}$ describes the spin relaxation process of ES with the damping rate $\Gamma_{\mathrm{JD}}$. Analytical expressions of $\mathcal{A}_{\mathrm{s}}^{\{\mathrm{ge}\}}$ , $\mathcal{A}_{\mathrm{q}}^{\{\mathrm{ge}\}}$ and $\mathcal{A}_{\mathrm{JD}}$ are derived in Refs. [1, 17]. The superoperator $\mathcal{A}_{\mathrm{p}}^{\{\mathrm{eg}\}}$ in Eq. (13) is defined as [9]
$R_{\mathrm{op}}\mathcal{A}_{\mathrm{p}}^{\{\mathrm{eg}\}}=$ $i\left(V^{*}{\otimes}W-W^{*}\otimes V\right),$ (14)
which represents the optical excitation of the atoms from the GS to the ES. Here, $\otimes$ denotes the Kronecker product between operators in Hilbert space.
III.2 Optical pumping in quasi-high-pressure regime
III.2.1 Depopulation
The energy denominator matrices $\Delta^{\{\mathrm{eg}\}}+i\Gamma_{\mathrm{brd}}$ in Eqs. (12) play a central role in describing optical pumping in different buffer-gas regimes. In the HP limit, the collisional broadening $\Gamma_{\mathrm{brd}}$ is much larger than the difference between the elements of the detuning matrix $\Delta^{\{\mathrm{eg}\}}$, i.e., $\Gamma_{\mathrm{brd}}\gg|\omega_{\bar{F},\bar{m}_{F}}^{\{\mathrm{e}\}}-\omega_{F,m_{F}}^{\{\mathrm{g}\}}|$. In this case, $\Delta^{\{\mathrm{eg}\}}$ is simplified to a single-valued matrix
$\Delta^{\{\mathrm{eg}\}}\approx\Delta_{0}\sum_{\bar{F},\bar{m}_{F},F,m_{F}}|\bar{F},\bar{m}_{F}\rangle\langle F,m_{F}|,$ (15)
and the effective Hamiltonian in Eq. (11) becomes
$\delta H^{\{\mathrm{g}\}}=\frac{V^{\dagger}V}{\Delta_{0}+i\Gamma_{\mathrm{brd}}}.$ (16)
Physically, the approximation in Eq. (15) means that, in the large broadening limit, the different detuning frequencies have a negligible effect on the pumping process, and the alkali-metal atom behaves like a simple two-level atom. With the symmetry property of the electric dipole operator $\hat{\bm{e}}^{*}\cdot\bm{d}\bm{d}^{\dagger}\cdot\hat{\bm{e}}=1/2-K\mathbf{s}\cdot\mathbf{S}$, the HP approximation Eqs. (15) & (16) allows a compact form of the depopulation pumping superoperator [9],
$\mathcal{A}^{(\mathrm{HP})}_{\mathrm{dp}}$= $1-K\mathbf{s}\cdot\left(\mathbf{S}^{\flat}+\mathbf{S}^{\sharp}\right)-\frac{i\Delta_{0}}{\Gamma_{\mathrm{brd}}}K\mathbf{s}\cdot\mathbf{S}^{\copyright},$ (17)
where $K=1$ corresponds to D1 line pumping, with K=1/2K=-1/2 for the D2 line, and $\mathbf{s}=i\hat{\bm{e}}\times\hat{\bm{e}}^{*}$ is the effective photon spin of the pump light. The real and imaginary parts of $\mathcal{A}_{\mathrm{dp}}^{(\mathrm{HP})}$ describe the light absorption and the optical shift, respectively. In Eq. (11), the optical pumping rate $R_{\mathrm{op}}=\bar{\sigma}\Phi$ is proportional to the incident photon flux $\Phi$, and $\bar{\sigma}$ is the intrinsic absorption cross-section of unpolarized atoms (i.e., all the GS levels are equally populated without pumping effect). The intrinsic absorption cross section in the HP limit is a function of detuning $\Delta_{0}$
$\bar{\sigma}_{\mathrm{HP}}=\frac{A}{\pi}\frac{\Gamma_{\mathrm{brd}}}{\Delta_{0}^{2}+\Gamma_{\mathrm{brd}}^{2}}\equiv A\mathcal{L}(\Delta_{0}),$ (18)
where $\mathcal{L}(\Delta)$ is the normalized Lorentzian lineshape function, and $A=\pi r_{\mathrm{e}}cf_{\mathrm{eg}}$ is the area constant with the classical electron radius $r_{\mathrm{e}}$, the speed of light $c$ and the atom oscillator strength $f^{\{ge\}}$.
In the QHP regime of interest, the ES hyperfine splittings and all the Larmor frequencies in weak magnetic fields are neglected, i.e., $\omega_{\bar{F},\bar{m}_{F}}^{\{\mathrm{e}\}}\approx 0$ and $\omega_{F,m_{F}}^{\{\mathrm{g}\}}\approx\omega_{F,0}^{\{\mathrm{g}\}}$, but the GS hyperfine splitting is kept. In this case, the elements of the detuning matrix $\Delta^{\{\mathrm{eg}\}}$ are grouped by two values $\Delta_{F}$ for $F=a$ or $b$, i.e.
$\Delta^{\{\mathrm{eg}\}}\approx\sum_{\bar{F},\bar{m}_{F},F,m_{F}}\Delta_{F}|\bar{F}\bar{m}_{F}\rangle\langle F,m_{F}|$ (19)
where $\Delta_{F}=\Delta_{0}+\omega^{\{\mathrm{g}\}}_{F,0}$ is the frequency detuning with respect to hyperfine split levels. With this coarse-grained modeling of the detuning matrix, the effective Hamiltonian $\delta H^{\{\mathrm{g}\}}$ becomes
$\delta H^{\{\mathrm{g}\}}=$ $\sum_{F=a,b}\delta H^{\{\mathrm{g}\}}_{F}\Pi_{F},$ (20)
$\delta H^{\{\mathrm{g}\}}_{F}$= $\frac{V^{\dagger}V}{\Delta_{F}+i\Gamma_{\mathrm{brd}}},$ (21)
where $\Pi_{F}=\sum_{m_{F}}|F,m_{F}\rangle\langle F,m_{F}|$ is the projection operator of the GS subspace with a given total angular momentum quantum number $F$, and $\delta H_{F}^{\{\mathrm{g}\}}$ is the $F$-resolved effective Hamiltonian.
With the QHP correction in Eq. (19), we show that the depopulation superoperator $\mathcal{A}_{\mathrm{dp}}$ becomes
$\mathcal{A}_{\mathrm{dp}}\mathcal{P}_{\mathrm{Z}}=\sum_{F}\left[1-K\mathbf{s}\cdot(\mathbf{S}^{\flat}+\mathbf{S}^{\sharp})-\frac{i\Delta_{F}}{\Gamma_{\mathrm{brd}}}K\mathbf{s}\cdot\mathbf{S}^{\copyright}\right]Q_{F}\mathcal{P}_{F}.$ (22)
See Appendix A for detailed proof. In comparison to the HP case Eq. (17), the hyperfine-resolved GS brings about the following corrections: (i) the overall detuning $\Delta_{0}$ is replaced by the $F$-resolved detuning $\Delta_{F}$; (ii) a weighting factor $Q_{F}$ is introduced as
$Q_{F}(\Delta_{F})=\frac{\mathcal{L}(\Delta_{F})}{x_{a}\mathcal{L}(\Delta_{a})+x_{b}\mathcal{L}(\Delta_{b})},$ (23)
where $x_{F}=(2F+1)/(2(2I+1))$ is the degeneracy weighting factor; and (iii) the expression in Eq. (22) is subjected in the Zeeman subspace with the projection operator in the Liouville space
$\mathcal{P}_{F}=$ $\sum_{m_{F},m^{\prime}_{F}}|F,m_{F};F,m^{\prime}_{F})(F,m_{F};F,m^{\prime}_{F}|,$ (24)
$\mathcal{P}_{\mathrm{Z}}=$ $\mathcal{P}_{a}+\mathcal{P}_{b}.$ (25)
The intrinsic absorption cross-section in the QHP regime should be modified as
$\bar{\sigma}_{\mathrm{QHP}}=A\left[x_{a}\mathcal{L}(\Delta_{a})+x_{b}\mathcal{L}(\Delta_{b})\right].$ (26)
Figure 2: Weighting factor $Q_{F}$ for $F=a$ and $F=b$ [see Eq. (23)], with the GS hyperfine splitting $6.8~\mathrm{GHz}$ and different values of collisional broadening $\Gamma_{\mathrm{brd}}=3,~10~\mathrm{and}~20\mathrm{~GHz}$. The curves converge to a common intersection at $\Delta_{a}=\Delta_{b}$, yielding $Q_{a}=Q_{b}=1$. The thick black dashed curve shows the maximum attainable $Q_{b}$ for increasing $\Gamma_{\mathrm{brd}}$, recovering the HP limit.
The corrected depopulation superoperator in Eq. (22) reveals the physical effect arising from the difference in absorption between hyperfine split levels. Indeed, this difference is characterized by the factor $Q_{F}$ with given $F=a$ or $F=b$, as shown in Fig. 2. For hyperfine coherence between levels $F=a$ and $F=b$, e.g., $|a,m_{a};b,m_{b})$, a simple form of $\mathcal{A}_{\mathrm{dp}}$ like Eq. (22) is not available, as proved in Appendix A. Fortunately, the hyperfine coherence is usually neglected in analyzing most of the observables like spin polarization, magnetic resonance linewidth, etc, as shown in Section IV.
III.2.2 Repopulation
The repopulation term in Eq. (13) represents the generation and relaxation of the atomic ES spin during the optical pumping process. In the similar way of dealing with $\mathcal{A}_{\mathrm{dp}}$, the optical excitation superoperator $\mathcal{A}^{\{\mathrm{eg}\}}_{\mathrm{p}}$ is simplified as (see Appendix A for detailed proof)
$\mathcal{A}^{\{\mathrm{eg}\}}_{\mathrm{p}}\mathcal{P}_{\mathrm{Z}}=2\sum_{F}(\mathbf{\hat{e}}^{*}\cdot\bm{d})\otimes(\bm{d}^{\dagger}\cdot\mathbf{\hat{e}})Q_{F}\mathcal{P}_{F}.$ (27)
This result physically means that detunings $\Delta_{F}$ for F=a,bF=a,\,b contribute differently to the optical excitation process.
For ES relaxation and ES-GS transition that occurs in Eq. (13), Appelt et al.. show that dominant mechanisms are typically J-damping collision and quenching [1]. Thus $\mathcal{A}_{\mathrm{rp}}$ can be expanded in power series of $H^{\{\mathrm{e}\}\copyright}$, retaining only the leading term
$\mathcal{A}_{\mathrm{rp}}\approx\mathcal{A}^{(0)}_{\mathrm{rp}}=\left(1-\mathcal{A}_{\mathrm{q}}^{\{\mathrm{ge}\}}\right)\left(1-\mathcal{A}_{\mathrm{JD}}\right)\mathcal{A}_{\mathrm{p}}^{\{\mathrm{eg}\}},$ (28)
while the first-order perturbation from $H^{\{\mathrm{e}\}\copyright}$ is given by
$\mathcal{A}^{(1)}_{\mathrm{rp}}=\left(1-\mathcal{A}_{\mathrm{q}}^{\{\mathrm{ge}\}}\right)\frac{iH^{\{\mathrm{ee}\}\odot}}{\Gamma_{\mathrm{q}}}\left(1-\mathcal{A}_{\mathrm{JD}}\right)\mathcal{A}_{\mathrm{p}}^{\{\mathrm{eg}\}}.$ (29)
Under lower pressure (e.g., $0.01\sim 0.1~\mathrm{amg}$ for $\mathrm{N_{2}}$), a systematic investigation in [17] offers a scale of correction from $H^{\{\mathrm{e}\}\copyright}$. For theoretical simplicity, we still adopt the approximation of Eq. (28) in the following analysis. As already researched [1, 9], J-Damping collsion and quenching fully transfer nuclear polarization with no transfer of electronic polarization. Furthermore, any mechanism originating from the electric-dipole interaction only manipulate the electronic part of a density matrix, and assuere nuclear spin invariant, due to its short interaction timescale. So superoperators $\mathcal{A}_{\mathrm{dp}}$ and $\mathcal{A}_{\mathrm{rp}}$ share a common nuclear part
$\mathcal{A}_{\mathrm{rp}}=\left(1-\mathcal{A}_{\mathrm{q}}^{\{\mathrm{ge}\}}\right)\mathcal{A}_{\mathrm{p}}^{\{\mathrm{eg}\}}=\left(1-\mathcal{A}_{\mathrm{SD}}\right)\mathcal{A}_{\mathrm{dp}},$ (30)
Figure 3: Real part of optical pumping superoperator within the Zeemnan subspace, for nuclear spin $I=3/2$, D line constant $K=1$, and mean photon spin $\mathbf{s}=\hat{z}$. The underlying matrix elements are presented in the HP limit and arranged in Liouville basis $|F,\,m_{F};F,\,m^{\prime}_{F})$. They are separated by dashed lines into $m$th- to $n$th- order Zeeman coherence blocks $\mathrm{Re}\left[\mathcal{A}_{\mathrm{op}}\right]_{m,n}$ (see Appendix. B). The block marked by the solid lines ($m=n=0$) governs population dynamics, while the block marked by the dotted lines ($m=n=1$) contributes to evolution of the 1st-order Zeeman coherence. The vertical gray and green bands encode $Q_{a}$ and $Q_{b}$ Eq. (23), acting as the QHP correction. The remaining blocks follow the same structure.
Through Eqs. (10), (22), and (30), we conclude that the full optical pumping superoperator $\mathcal{A}_{\mathrm{op}}$ with the QHP correction is
$\mathcal{A}_{\mathrm{op}}\mathcal{P}_{\mathrm{Z}}=$ $\sum_{F}\left[\mathcal{A}_{\mathrm{SD}}-\frac{K\mathbf{s}}{2}\cdot\bm{\mathcal{A}}_{\mathrm{SE}}-\frac{i\Delta_{F}}{\Gamma_{\mathrm{brd}}}K\mathbf{s}\cdot\mathbf{S}^{\copyright}\right]Q_{F}\mathcal{P}_{F}.$ (31)
The compact form containing the $F$-resolved counterpart enables us to revisit the master equation Eq. (6) in the QHP regime. For higher computational efficiency and clearer physical interpretation in Liouville space, matrix elements of the superoperators are grouped into blocks according to the characteristic evolution frequencies of the density-matrix elements. The technical details of this procedure are provided in the Appendix B. Figure 3 illustrates the matrix elements and the corresponding block structure of $\mathrm{Re}\left[\mathcal{P}_{\mathrm{Z}}\mathcal{A}_{\mathrm{op}}\mathcal{P}_{\mathrm{Z}}\right]$, which represents the QHP-corrected optical pumping superoperator restricted to the Zeeman subspace of primary interest.
III.3 Observables in quasi-high-pressure regime
In this subsection, we investigate observables in the QHP regime for the common configuration in which the applied magnetic field is collinear with the pump beam. Specifically, we analyze the optical absorption, population distributions, spin polarization, and magnetic-resonance linewidth, and we highlight the ways in which their behavior departs from that observed in the HP regime.
III.3.1 Light absorption of atoms
Optical absorption of atoms is quantified by the absorption cross section $\sigma$, which is expressed in terms of the effective Hamiltonian $\delta H^{\{\mathrm{g}\}}$ and the GS population $|\rho_{0})$ [9].
$\sigma$ $=-\dfrac{2}{\Phi}\left(\mathrm{Im}\left.\left[\delta H^{\{\mathrm{g}\}}\right]\right|\rho_{0}\right),$ (32)
where $(A|B)=\mathrm{Tr}[AB]$ denotes the inner product in Liouville space. Following Eqs. (20), (21) and (26), the absorption cross-section in the QHP regime is
$\sigma$ $=A\left[C_{a}\mathcal{L}(\Delta_{a})+C_{b}\mathcal{L}(\Delta_{b})\right].$ (33)
with the weighting factor $C_{F}$ defined by
$C_{F}$ $=p_{F}-{2Ks\langle S_{z}\rangle_{F}},$ (34)
where $s$ is effective photon spin along the $\hat{z}-$axis ($s=0$ or $\pm 1$ stands for the linearly or circularly polarized beam), and $p_{F}$ and $\langle S_{z}\rangle_{F}$ are the total population and the spin polarization of the states with a given $F$, i.e.
$p_{F}=$ $\sum_{m_{F}}\langle F,m_{F}|\rho_{0}|F,m_{F}\rangle,$ (35)
$\langle S_{z}\rangle_{F}=$ $\sum_{m_{F}}\langle F,m_{F}|S_{z}|F,m_{F}\rangle\langle F,m_{F}|\rho_{0}|F,m_{F}\rangle.$ (36)
The $C_{F}$ coefficient Eq. (34) comprises two terms representing distinct optical pumping processes. The $F$-resolved population term, $p_{F}$, characterizes the population redistribution between the $F=a$ and $F=b$ levels. Meanwhile, the $\langle S_{z}\rangle_{F}$ term accounts for population transfer among the $m_{F}$ sublevels, reflecting the generation of spin polarization through angular momentum exchange with incident photons. In the following of this paper, these two mechanisms are referred to as the $F$-pumping process and $m_{F}$-pumping process, respectively.
In the HP limit, the two Lorentzian lineshapes exhibit degeneracy, such that $\mathcal{L}_{a}(\Delta_{a})\approx\mathcal{L}_{b}(\Delta_{b})\approx\mathcal{L}(\Delta_{0})$, thereby suppressing the $F$-pumping process. Consequently, the $C_{F}$ coefficient reduces to the form $C=1-2Ks\langle S_{z}\rangle$, indicating that absorption is predominantly governed by spin polarization driven by $m_{F}$-pumping. In contrast, within the QHP regime, both $F$-pumping and $m_{F}$-pumping processes are integral to the system dynamics. Therefore, the $F$-resolved population, $p_{F}$, and the spin polarization, $\langle S_{z}\rangle_{F}$, encapsulate the fundamental mechanisms governing the QHP regime. Elucidating the dynamics of these parameters as a function of pump power and light frequency is crucial for uncovering the underlying principles of optical pumping.
III.3.2 Population and polarization
In the presence of longitudinal pumping light, spin-destruction and spin-exchange relaxation in Eq. (6), the population dynamics are governed by
$\frac{d\left|\rho_{0}\right)}{dt}=-\mathcal{G}_{\mathrm{pop}}\left|\rho_{0}\right).$ (37)
with the superoperator $\mathcal{G}_{\mathrm{pop}}$
$\mathcal{G}_{\mathrm{pop}}=$ $\Gamma_{\mathrm{SD}}\left[\mathcal{A}_{\mathrm{SD}}\right]_{0,0}+R_{\mathrm{op}}\left[\mathcal{A}_{\mathrm{op}}\right]_{0,0}$
$+\Gamma_{\mathrm{ex}}\left(\left[\mathcal{A}_{\mathrm{SD}}\right]_{0,0}-\langle S_{z}\rangle\left[\mathcal{A}^{(z)}_{\mathrm{SE}}\right]_{0,0}\right),$ (38)
where $\langle S_{z}\rangle=(S_{z}|\rho_{0})$ is the expectation value of the electron spin along the $z$ direction.
The steady-state solution, $\left|\rho_{\mathrm{ss}}\right)=\left|\rho_{0}(t\to+\infty)\right)$, is obtained by evaluating the null space of $\mathcal{G}_{\mathrm{pop}}$. Within the QHP regime, the population ratio, $\eta$, between the $|F=a,m_{F}\rangle$ and $|F=b,m_{F}\rangle$ states for a given $m_{F}$ ($|m_{F}|\neq a$) is given by
$\eta=\frac{(a,m_{F};a,m_{F}\left|\rho_{\mathrm{ss}}\right)}{(b,m_{F};b,m_{F}\left|\rho_{\mathrm{ss}}\right)}=\frac{\Gamma_{\mathrm{ex}}+\Gamma_{\mathrm{SD}}+R_{\mathrm{op}}{Q}_{b}}{\Gamma_{\mathrm{ex}}+\Gamma_{\mathrm{SD}}+R_{\mathrm{op}}{Q}_{a}}.$ (39)
This ratio is independent of $m_{F}$. In the conventional HP limit, the absorption efficiencies become equivalent ($Q_{a}=Q_{b}$), driving the ratio to $\eta\rightarrow 1$. This limit seamlessly recovers the widely recognized spin-temperature distribution (STD) [11, 1]. However, in the scenario where $Q_{a}\neq Q_{b}$, the system diverges from this standard behavior. Under these conditions, the asymmetric absorption efficiencies parameterized by $Q_{F}$ induce an effective population transfer between the $F=a$ and $F=b$ multiplets, a phenomenon that dictates the core dynamics explored in the subsequent analysis.
For linearly polarized pumping with effective photon spin $s=0$, the population distribution is exactly solvable. Since no angular momentum is transferred from the incident field to the atomic system, the $m_{F}$-pumping process is precluding. Consequently, these magnetic sublevels remain uniformly populated, and the generation of spin polarization is strictly forbidden (i.e., $\langle S_{z}\rangle_{F}=\langle S_{z}\rangle\equiv 0$). The system dynamics therefore reduce to those of an effective two-level system comprising the hyperfine multiplets $F=a$ and $F=b$. Under these conditions, the analytical steady-state solution, $\left|\rho_{\mathrm{lin}}\right)$, dictated by the principle of detailed balance [16], is given by
$\left|\rho_{\mathrm{lin}}\right)$ $=\sum_{F,m_{F}}\frac{p_{F}}{2F+1}\left|F,m_{F};F,m_{F}\right),$ (40)
$p_{F}$ $=\frac{x_{F}\left(\Gamma_{\mathrm{SD}}+\Gamma_{\mathrm{ex}}+R_{\mathrm{op}}Q_{F^{\prime}\neq F}\right)}{\Gamma_{\mathrm{SD}}+\Gamma_{\mathrm{ex}}+R_{\mathrm{op}}(x_{a}Q_{b}+x_{b}Q_{a})},$ (41)
where $\Gamma_{\mathrm{ex}}+\Gamma_{\mathrm{SD}}$ denotes the relaxation-induced population exchange rate between the multiplets, acting as a restorative mechanism that drives the system toward the equilibrium. Conversely, the term $R_{\mathrm{op}}Q_{F^{\prime}\neq F}$ characterizes the effective population transfer from the $F^{\prime}$ to the $F$ multiplet, creating the unbalanced population between the two manifolds under the linearly polarized pumping.
Figure 4: (a) Representative steady-state populations under slow spin-exchange conditions.. (b) Total population of the $F=a$ multiplet as a function of the normalized pumping rate $R_{\mathrm{op}}/\Gamma_{\mathrm{SD}}$, with curves denoting different $Q_{F}$ configurations. (c) The 1-norm distance $d$ [see Eq. (42)] between the steady-state population and an STD profile of the equivalent polarization, fixing $R_{\mathrm{op}}/\Gamma_{\mathrm{SD}}=1$. (d) Spin polarization $P$ versus $\Gamma_{\mathrm{ex}}/\Gamma_{\mathrm{SD}}$. Color coding for (c) and (d) is consistent with (b). Results are obtained for $I=3/2$ and $K=s=1$.
Under circularly polarized pumping ($s=\pm 1$), the system dynamics become increasingly complex due to the concurrent action of $F$-pumping and $m_{F}$-pumping. Because these coupled processes are governed by the non-linear Eq. (37), which lacks an analytical solution, the steady-state distributions are solved via numerical simulation.
Specifically, in the standard HP limit, solely $m_{F}$-pumping dictates the dynamics; this mechanism drives the population toward states with higher $m_{F}$ values while maintaining the population balance between the $F=a$ and $F=b$ manifolds for a given $m_{F}$. Conversely, within the QHP regime, the emergence of $F$-pumping breaks this multiplet balance. As illustrated in Fig. 4(a), when $Q_{a}>Q_{b}$, the disproportionate absorption within the $F=a$ manifold pumps the population into the $F=b$ levels (red bars). An analogous population redistribution occurs under the $Q_{b}>Q_{a}$ condition (blue bars). Figure 4(b) elucidates the dependence of this impact on the normalized pumping intensity, parameterized as $R_{\mathrm{op}}/\Gamma_{\mathrm{SD}}$. This redistribution mechanism exerts a negligible influence when approaching either the weak-pumping limit ($R_{\mathrm{op}}\ll\Gamma_{\mathrm{SD}}$) of a nearly uniform distribution or the strong-pumping limit ($R_{\mathrm{op}}\gg\Gamma_{\mathrm{SD}}$) of stretched-state saturation. In contrast, the intermediate domain obviously manifests the $F$-pumping dynamics driven by $Q_{a}\neq Q_{b}$, where, for instance, the notably suppressed red curve (compared with the dashed line) illustrates the net transfer from the $F=a$ to the $F=b$ levels.
The $F$-pumping-induced imbalance in the QHP regime drives the system away from the conventional STD characteristic of the HP limit. Specifically, the steady-state density matrix can no longer be parameterized by a single temperature factor $\beta(P)$ in the form $|\rho_{\mathrm{STD}})\propto\exp[\beta(P)F_{z}]$ [11, 1], with $F_{z}=S_{z}+I_{z}$ being the $z$-component of the total angular momentum operator ($I_{z}$ the nuclear spin counterpart), and $P=2(S_{z}|\rho_{\mathrm{ss}})$ being the steady-state spin polarization . To quantify this deviation, we define a 1-norm distance metric
$d=\|\rho_{\mathrm{ss}}-\rho_{\mathrm{STD}}\|_{1}$ (42)
which compares the QHP steady-state distribution to an ideal STD possessing identical spin polarization $P$. Figure. 4(c) plots this distance as a function of $\Gamma_{\mathrm{ex}}/\Gamma_{\mathrm{SD}}$—the ratio of the spin-exchange rate to the spin-destruction rate—at a fixed optical pumping rate $R_{\mathrm{op}}$. The distance $d$ decreases monotonically as the ratio increases. This demonstrates that rapid spin-exchange collisions provide a dominant restorative mechanism, effectively thermalizing the system to counteract the $F$-pumping induced by asymmetric absorption. The high validity of the spin-temperature approximation under the strong spin-exchange condition ($\Gamma_{\mathrm{ex}}/\Gamma_{\mathrm{SD}}\gg 1$) establishes a critical theoretical foundation for utilizing the STD to simplify the physical models of various practical applications, such as SERF magnetometers, even when the HP condition is not necessarily fully satisfied.
Although the STD serves as a valid approximation at high spin-exchange rate, the joint contribution of $F$- and $m_{F}$-pumping under QHP condition gives rise to a spin polarization different to that obtain in the HP condition. In the HP limit with $Q_{a}=Q_{b}$ [see the black dashed line in Fig. 4(c)], the spin polarization is
$P_{\mathrm{HP}}=\frac{KsR_{\mathrm{op}}}{R_{\mathrm{op}}+\Gamma_{\mathrm{SD}}}$ (43)
In contrast, when $Q_{a}\neq Q_{b}$, the polarization $P_{\mathrm{QHP}}$ converges to a stable value at high spin-exchange rate. The asymptotic polarization under QHP conditions (blue and red dashed lines) is sensitive to the asymmetric absorption strengths ($Q_{a}$ and $Q_{b}$). Consequently, within the QHP framework, the spin polarization $P$ under the STD approximation cannot be obtained trivially; and must be reshaped by the imbalance $Q_{a}\neq Q_{b}$. Detailed results and their corresponding observable effects will be presented in Section IV.
III.3.3 Intrinsic magnetic resonance
The population imbalance across the $m_{F}$ sublevels enables Zeeman magnetic resonance under a radio-frequency (RF) driving field oscillating at $\omega_{\mathrm{rf}}$. To streamline the theoretical analysis, we conceptualize the framework established in our previous work [22] under the notion of intrinsic magnetic resonance. This intrinsic regime describes the 1st-order Zeeman coherence generated by an infinitesimally weak driving field, operating under a weak static magnetic field where the non-linear Zeeman effect is negligible, so that the spin precession frequencies degenerate into a singular Larmor frequency $\omega_{\mathrm{L}}$. In this case, the resonance dynamics are captured by the generated 1st-order Zeeman coherence $\left|\tilde{\rho}_{1}\right)$. In the RF rotating frame, the coherencce is governed by $\left|\tilde{\rho}_{1}\right)\sim\Omega_{\mathrm{R}}\mathcal{G}^{-1}_{\mathrm{\mathrm{WDA}}}\big[S^{\copyright}_{x}\big]_{1,0}\left|{\rho}_{\mathrm{st}}\right),$ where $\Omega_{\mathrm{R}}$ is the weak Rabi frequency, $\big[S^{\copyright}_{x}\big]_{1,0}$ dictates the coherence generation from polarized atoms, and $\mathcal{G}_{\mathrm{WDA}}$ is the total relaxation superoperator under the weak driving approximation. When the resonance is measured by observing the transverse spin component via Faraday rotation [10, 4], and the signal is characterized by the complex amplitude
$R=(S_{x}\left|\tilde{\rho}_{1}\right)\approx$ $\frac{\Omega_{\mathrm{R}}}{2}\frac{W_{\mathrm{}}}{\Delta_{\mathrm{rf}}+i\Gamma_{2}}$ (44)
Here, $W$ represents the dominant Zeeman coherence weight, and $\Delta_{\mathrm{rf}}=\omega_{\mathrm{rf}}-\omega_{\mathrm{L}}-\delta\omega_{\mathrm{L}}$ defines the effective RF detuning, taking into account the optical shift $\delta\omega_{\mathrm{L}}$. The linewidth $\Gamma_{2}$ is obtained by the minimal eigenvalue of $\mathcal{G}_{\mathrm{WDA}}$, and the optical shift originates from the value of $\mathrm{Im}[\mathcal{A}_{\mathrm{op}}]_{1,1}$ [22]. Experimentally, properties of the resonance directly manifest as the in-phase and out-of-phase quadratures, i.e., $X=\mathrm{Re}[R]$ and $Y=-\mathrm{Im}[R]$, respectively.
Transitioning this framework into the QHP regime necessitates a $Q_{F}$-induced modification. While all other relaxation mechanisms in $\mathcal{G}_{\mathrm{WDA}}$ remain formally unchanged, the optical pumping contribution to $\left|\tilde{\rho}_{1}\right)$ relaxation is replaced by the hyperfine-resolved superoperator $\left[\mathcal{A}_{\mathrm{op}}\right]_{1,1}$ [see Eq. (31) & Fig. 3] to incorporate the $Q_{a}\neq Q_{b}$ absorption asymmetry. Detailed results about the linewidth calculations and the magnetic resonance enhancement under QHP conditions are discussed in Sec. IV.3.
IV RESULTS AND DISCUSSION
IV.1 Absorption cross section
For optically pumped atoms in the QHP regime, we utilize the description of light absorption in Sec. III.3.1 to calculate the absorption cross-section $\sigma_{\mathrm{lin}}$ and $\sigma_{\mathrm{cir}}$ for linearly- and circularly-polarized laser beams.
For linearly polarized pumping ($s=0$), populations in the hyperfine multiplets $F=a$ and $F=b$ contribute to the light absorption differently, and $\sigma_{\mathrm{lin}}$ is
$\sigma_{\mathrm{lin}}$ $=A\left[p_{a}\mathcal{L}(\Delta_{a})+p_{b}\mathcal{L}(\Delta_{b})\right],$ (45)
where the $F$-multiplet populations $p_{a}$ and $p_{b}$ are obtained in Eq. (41). The results of $\sigma_{\mathrm{lin}}$ are shown in Fig. 5(a), where a notable feature emerges for larger $\Phi$. In the QHP regime, the resonance peaks of $\sigma_{\mathrm{lin}}$ for pumped atoms become significantly suppressed and unresolved. This is evident in Fig. 5(a), while it is absent in the HP limit as shown in Fig. 5(b). The behavior can be understood from the fact that, when pumping near $\Delta_{F}=0$, population in a given hyperfine manifold $F$ is more efficiently pumped into the other manifold $F^{\prime}\neq F$, which leads to a reduction of the absorption. As the photon flux $\Phi$ continues to increase, this $F$-pumping mechanism is further enhanced, eventually driving $p_{F}$ almost entirely into a single hyperfine multiplet. Consequently, in the strong-pumping limit $R_{\mathrm{op}}\gg\Gamma_{\mathrm{SD}}+\Gamma_{\mathrm{ex}}$ (assuming negligible population trapping in the ES), the cross section asymptotically approaches a finite functional form:
$\sigma_{\mathrm{lin}}(R_{\mathrm{op}}\gg\Gamma_{\mathrm{SD}}+\Gamma_{\mathrm{ex}})$ $=\frac{A_{\mathrm{eff}}}{\pi}\frac{\Gamma_{\mathrm{eff}}}{\Delta^{2}_{0}+\Gamma^{2}_{\mathrm{eff}}},$ (46)
which exhibits a single-Lorentzian profile centered at the reference frequency, i.e., $\Delta_{0}=x_{a}\omega^{\{\mathrm{g}\}}_{a,0}+x_{b}\omega^{\{\mathrm{g}\}}_{b,0}=0$. Here, the effective spectral area and linewidth are $A_{\mathrm{eff}}=A/\xi$ and $\Gamma_{\mathrm{eff}}=\xi\,\Gamma_{\mathrm{brd}}$, respectively. The ratio $\xi$ is
$\xi$ $=\sqrt{1+x_{a}x_{b}\left(\dfrac{\omega^{\{\mathrm{g}\}}_{\mathrm{hf}}}{\Gamma_{\mathrm{brd}}}\right)^{2}},$ (47)
[…]
Throughout the process of varying $\Phi$, a notable feature is the emergence of a unique stationary point at the centered detuning
[…]
where $\sigma_{\mathrm{lin}}$ remains independent of $\Phi$. This characteristic point arises because the light is equally detuned from both multiplets ($\Delta_{a}=\Delta_{b}$). Such equal absorption strength effectively switches off the $F$-pumping dynamics, forcing the cross section to remain fixed at $\bar{\sigma}_{\mathrm{QHP}}$ Eq. (26).
Figure 5: Absorption cross section $\sigma$ versus light detuning. (a)–(b) Results for linear polarized pumping ($s=0$), under typical QHP ($2\Gamma_{\mathrm{brd}}=3~\mathrm{GHz}$) and HP ($2\Gamma_{\mathrm{brd}}=20~\mathrm{GHz}$) conditions. (c)–(d) Analogous plots for circular polarized pumping ($s=\pm 1$) on the $D_{1}$ line. In each panel, the vertical dotted lines denote the main resonance centers. The solid curves, arranged from top to bottom, span a photon flux $\Phi$ range from 0 to $10^{18}~\mathrm{cm^{-2}\cdot s^{-1}}$ for $s=0$, and from 0 to $5\times 10^{14}~\mathrm{cm^{-2}\cdot s^{-1}}$ for $s=\pm 1$. The 87Rb atoms are considered for simulations, with a fixed $\Gamma_{\mathrm{ex}}=7.75~\mathrm{kHz}$. To account for the spin-destruction collision induced by buffer gas, $\Gamma_{\mathrm{SD}}$ is set to $0.05~\mathrm{kHz}$ and $0.27~\mathrm{kHz}$ for the QHP and HP cases.
For circularly polarized pumping ($s=\pm 1$), we focus on the commonly used $D_{1}$ transition ($K=1$). Using the STD approximation detailed in Sec. III.3.2, we substitute the expressions of $p_{F}$ and $\langle S_{z}\rangle_{F}$ from Table 1 into (34). Consequently, the absorption cross section $\sigma_{\mathrm{cir}}$ is expressed as
$\sigma_{\mathrm{cir}}=A\left[C_{a}(I,\,P)\mathcal{L}(\Delta_{a})+C_{b}(I,\,P)\mathcal{L}(\Delta_{b})\right].$ (49)
where $C_{F}$ becomes an explicit function of nuclear spin $I$ and the polarization $P$. Taking $I=3/2$ as an example, the coefficients in Eq. (34) are
$C_{a=2}$ $=\frac{5}{8}-\frac{10P-5P^{2}+6P^{3}-P^{4}}{8(1+P^{2})},$ (50)
$C_{b=1}$ $=\frac{3}{8}+\frac{2P-5P^{2}-2P^{3}-P^{4}}{8(1+P^{2})},$ (51)
Representative spectra under QHP and HP conditions are displayed in Figs. 5(c) and 5(d), where an overall reduction in absorption as $\Phi$ increases is common to the both cases, mainly driven by the population accumulated in the dark stretched state $|F=a,m_{F}=sa\rangle$. Notably, Figure 5(c) shows an observation featured in the QHP regime: the higher resonance peak shifts from $\Delta_{a}=0$ at low $\Phi$ to $\Delta_{b}=0$ at high $\Phi$. The threshold photon flux equalizing these two peaks is denoted by $\Phi_{\mathrm{eq}}$ (solid black line). This inversion originates from two competing mechanisms. At low $\Phi$ values, low-polarization absorption is governed by the higher degeneracy of the $F=a$ multiplet (more available $m_{F}$ levels). As polarization builds up, the dominant absorbing states becomes $|F,m_{F}=s|a-1|\rangle$, and the state for $F=b$ contains a larger fraction of the electronic spin component capable of absorbing photons (e.g., spin-down electrons for $s=1$). The critical point of equal-absorption at $\Phi_{\mathrm{eq}}$ occurs when these two competing effects are balanced at $\Delta_{F}=0$ and lead to $C_{a}=C_{b}$. The corresponding polarization behavior under this condition is analyzed in the following subsection.
Table 1: Analytic expressions for $p_{a}$ and $\langle S_{z}\rangle_{a}$ defined in Eqs. (35) & (36), These relations are derived under the STD approximation, as a function of nuclear spin $I$ and the spin polarization $P$. The $F=b$ counterparts follow as $p_{b}=1-p_{a}$ and $\langle S_{z}\rangle_{b}=P/2-\langle S_{z}\rangle_{a}$. The rightmost column lists the equal-polarization point $P_{\mathrm{eq}}$ for D1 line pumping, obtained by solving Eq. (57) with $K=|s|=1$. For $I>1/2$, these results exclude the trivial $P=1$ root.
II $p_{a}(I,P)$ $\langle S_{z}\rangle_{a}(I,P)$ $|P_{\mathrm{eq}}|$
$\dfrac{1}{2}$ $\dfrac{3}{4}+\dfrac{P^{2}}{4}$ $\dfrac{P}{2}$ 1
11 $\dfrac{2}{3}+\dfrac{4P^{2}}{3(3+P^{2})}$ $\dfrac{5P+P^{3}}{3(3+P^{2})}$ 0.394
$\dfrac{3}{2}$ $\dfrac{5}{8}+\dfrac{5P^{2}+P^{4}}{8(1+P^{2})}$ $\dfrac{5P+3P^{3}}{8\left(1+P^{2}\right)}$ 0.207
22 $\dfrac{3}{5}+\dfrac{4(5P^{2}+3P^{4})}{5(5+10P^{2}+P^{4})}$ $\dfrac{3P+10P^{3}+3P^{5}}{2(5+10P^{2}+P^{4})}$ 0.128
$\dfrac{5}{2}$ $\dfrac{7}{12}+\dfrac{35P^{2}+42P^{4}+3P^{6}}{12\left(3+10P^{2}+3P^{4}\right)}$ $\dfrac{7P+14P^{3}+3P^{5}}{3\left(3+10P^{2}+3P^{4}\right)}$ 0.087
Beyond the specific spectral profiles against light detuning, the total light-absorption strength can be characterized by the integrated spectral area
$\alpha=\int\sigma(\Delta_{0})\,d\Delta_{0}.$ (52)
Figure 6 compares the variation of $\alpha$ for different $s$ in both QHP and HP cases. For linearly polarized light (black curves), $\alpha$ decays from $A$ and plateaus at $A_{\mathrm{eff}}$ in the strong-pumping limit, consistent with Eq. (46). Notably, the QHP case exhibits a pronounced reduction in $\alpha$ (solid line), whereas the HP case remains nearly unchanged (dashed line). This behavior is dictated by the ratio $\xi$ defined in Eq. (47)]: the HP limit $\Gamma_{\mathrm{brd}}\gg\omega^{\{\mathrm{g}\}}_{\mathrm{hf}}$ keeps $\xi\approx 1$ and suppresses $F$-pumping, whereas the QHP regime $\Gamma_{\mathrm{brd}}\sim\omega^{\{\mathrm{g}\}}_{\mathrm{hf}}$ allows active $F$-pumping to deplete absorption ($\xi>1$). Conversely, circular pumping (red curves) drives $\alpha$ strictly to zero, as $m_{F}$-pumping perfectly polarizes atoms into the non-absorbing stretched state. This vanishing is faster under QHP conditions due to stronger resonant absorption.
Figure 6: Integrated spectral area $\alpha$ [see Eq. (52)] normalized by AA versus photon flux $\Phi$, incorporating both linear ($s=0$) and circular ($s=\pm 1$, $D_{1}$ line) polarized pumping. Solid and dashed lines correspond to the QHP and HP cases, respectively. With increasing $\Phi$, the $s=0$ curves asymptotically converge to $\xi=0.41$ and $0.95$ [see Eq. (47)]. Parameters are consistent with Fig. 5.
IV.2 Spin polarization
Building upon the population dynamics in Sec. III.3.2, we study the steady-state spin polarization $P$ in the QHP regime. To start with, applying $(F_{z}|=(I_{z}|+(S_{z}|$ to both sides of Eq. (37) yields the evolution equation for the average total angular momentum $\langle F_{z}\rangle$:
$\frac{d\langle F_{z}\rangle}{dt}=$ $2R_{\mathrm{op}}\sum_{F}\mathcal{Q}_{F}\left(Ksp_{F}-2\langle S_{z}\rangle_{F}\right)-\Gamma_{\mathrm{SD}}\langle S_{z}\rangle,$ (53)
In obtaining this equation, we apply the superoperator identities in the population subspace
[…]
by which the nonlinear $\Gamma_{\mathrm{ex}}$ term vanishes in the $\langle F_{z}\rangle$ evolution, and meanwhile, the optical pumping superoperator $\left[A_{\mathrm{op}}\right]_{0,0}$ contributes the $p_{F}$ and $\langle S_{z}\rangle_{F}$ terms. They effectively acts as $F$-resolved spin-exchange and spin-destruction processes, which respectively capture the distinct impacts of $F$-pumping and $m_{F}$-pumping on the overall polarization dynamics.
Figure 7: (a) Two–dimensional color map of spin polarization $P$, with $K=|s|=1$ and $2\Gamma_{\mathrm{brd}}=3$ GHz. The white line signifies the photon flux $\Phi^{\prime}_{\mathrm{eq}}=7.6\times 10^{12}~\mathrm{cm^{-2}\cdot s^{-1}}$ that leads to $P(\Delta_{a}=0)=P(\Delta_{b}=0)=P_{\mathrm{eq}}$. (b) Variation of polarization with light detuning at representative $\Phi_{\mathrm{}}$ values. The black dashed line isolates the $\Phi^{\prime}_{\mathrm{eq}}$ case, alongside the solid red and blue curves for $\Phi=2\times 10^{12}$ and $2\times 10^{13}~\mathrm{cm^{-2}\cdot s^{-1}}$. The spin-destruction rate for 87Rb atoms used here is fixed at $\Gamma_{\mathrm{SD}}=0.05~\mathrm{kHz}$.
By setting $d\langle F_{z}\rangle/dt=0$, we obtain the implicit steady-state equation for the polarization:
$P=\frac{KsR_{\mathrm{op}}({Q}_{a}p_{a}+{Q}_{b}p_{b})}{\displaystyle R_{\mathrm{op}}\left({Q}_{a}u_{a}+{Q}_{b}u_{b}\right)+\Gamma_{\mathrm{SD}}},$ (56)
where $u_{F}=\left\langle S_{z}\right\rangle_{F}/\left\langle S_{z}\right\rangle$ represents the polarization weight of the $F$ multiplet. Adopting the STD approximation, we utilize the explicit expressions in Table 1 to convert Eq. (56) into a polynomial equation of $P$, which can be readily solved once the rate parameters are given. Under typical QHP conditions, the resulting $P$ is shown in Fig. 7, revealing several distinct features. Figure 7(a) demonstrates that $P$ generally builds up with increasing $\Phi$, and manifests two distinct maxima at $\Delta_{a}=0$ and $\Delta_{b}=0$ when the hyperfine structure are well-resolved. Figure 7(b) profiles several representative slices, highlights a shifting dominance across distinct $\Phi$ levels: lower photon fluxes favor polarization via the $\Delta_{a}=0$ resonance (red line), whereas higher fluxes make $\Delta_{b}=0$ more efficient (blue line). This behavior mirrors the $D_{1}$ absorption cross section $\sigma_{\mathrm{cir}}$ disscussed in Sec. IV.1; stronger resonant absorption directly dictates a higher rate of angular momentum injection, thus yielding a higher $P$.
Importantly, an equal-polarization point emerges at a critical flux $\Phi=\Phi^{\prime}_{\mathrm{eq}}$, where the polarization values at the $F=a$ and $F=b$ resonances equalize at a common value $P_{\mathrm{eq}}$. This phenomenon is denoted by the white dashed line in Fig. 7(a) and the black curve in Fig. 7(b). From Eq. (56), solving for the $P_{\mathrm{eq}}$ critical condition is equivalent to
$Ksp_{a}-2\langle S_{z}\rangle_{a}$ $=Ksp_{b}-2\langle S_{z}\rangle_{b},$ (57)
with the corresponding flux calculated as
$\Phi^{\prime}_{\mathrm{eq}}$ $=\frac{P_{\mathrm{eq}}\Gamma_{\mathrm{SD}}}{A(Ksp_{F}-2\langle S_{z}\rangle_{F})[\mathcal{L}(0)+\mathcal{L}(\omega^{\{\mathrm{g}\}}_{\mathrm{hf}})]}$. (58)
Here, Equation (57) signifies a fundamental balance independent of collisional broadening or any relaxation rate. This physically indicates that the combined optical pumping effects (including $F$- and $m_{F}$-pumping) exert an identical impact on $P$ through $F=a$ and $F=b$ levels. Under the STD approximation, Eq. (57) becomes a polynomial equation of $P_{\mathrm{eq}}$ uniquely determined by the nuclear spin $I$ (as listed in Table 1). We observe that a larger $I$ leads to a higher $P_{\mathrm{eq}}$. This stems from the fact that a larger $I$ creates more $m_{F}$ sublevels, which dilutes the initial low-$\Phi$ degeneracy advantage of $F=a$, allowing the $F=b$ resonance to rapidly overtake it in polarization efficiency.
Finally, it is worth noting that for the $D_{1}$ line ($K=1$), this equal-polarization state is exactly identical to the equal-absorption point $(C_{a}=C_{b})$, obtained by the common photon flux $\Phi_{\mathrm{eq}}=\Phi^{\prime}_{\mathrm{eq}}$. However, this correspondence does not hold for the $D_{2}$ line (K=1/2K=-1/2) due to its fundamentally different optical pumping pathways.
Figure 8: (a) Light-induced contribution to the magnetic resonance linewidth, $\delta\Gamma_{2}$, evaluated at D1 line optical resonance. The solid and dashed curves denote the QHP and HP cases. The shaded regions highlight the narrowing effect when the pumping light targets the $F=b$ level in the QHP regime. (b) and (c) display the normalized in-phase and out-of-phase quadratures ($X$ and $Y$) of the complex response $R$ [see Eq. (44)]. The red and blue curves correspond to the colored markers in (a), representing $\Delta_{a}=0$ and $\Delta_{b}=0$. For these calculations, we consider 87Rb atoms subjected to a weak magnetic field $B=5~\mu\mathrm{T}$, with detailed relaxation parameters matching those in figures in Sec. V.
IV.3 Magnetic resonance enhancement
For polarized atoms under RF driving, we specifically investigate the impact of QHP-specific optical pumping on $\Gamma_{2}$, which governs the performance of the intrinsic magnetic resonance discussed in Sec. III.3.3. This impact is explicitly described by the light-induced linewidth variation
$\delta\Gamma_{2}=\Gamma_{2}-\Gamma_{2,\,\mathrm{col}}.$ (59)
Here, $\Gamma_{2,\,\mathrm{col}}$ serves as the baseline linewidth obtained at vanishing photon flux i.e, $P\rightarrow 0$. For nuclear spin $I$, it takes the form determined by the spin-destruction and spin-exchange collision rates
$\Gamma_{2,\,\mathrm{col}}=$ $q_{\mathrm{SD}}(I)\Gamma_{\mathrm{SD}}+q_{\mathrm{ex}}(I)\Gamma_{\mathrm{ex}},$ (60)
with $q_{\mathrm{SD}}(I)$ and $q_{\mathrm{ex}}(I)$ the slowing-down factors [22]
$q_{\mathrm{SD}}(I)=\frac{(2I+1)^{2}}{2I^{2}+I+1},$ (61)
$q_{\mathrm{ex}}(I)=\frac{3(2I+1)^{2}}{2I(2I-1)}.$ (62)
The light-independent form of $\Gamma_{2,\,\mathrm{col}}$ remains invariant to buffer gas pressure, which only linearly scales $\Gamma_{\mathrm{SD}}$.
For representative QHP and HP conditions, Figure 8(a) plots $\delta\Gamma_{2}$ at optical resonance against the photon flux $\Phi$. For the traditional HP case (dashed line), the $F$-unresolved behavior of $\delta\Gamma_{2}$ is expected. Under rapid spin-exchange conditions, it initially decreases and then increases, reproducing the classical linewidth narrowing and broadening phenomena in specific polarization ranges [22, 2]. While $\delta\Gamma_{2}$ in the QHP regime (solid lines) preserves this behavior with increasing $\Phi$, it manifests pronounced $F$-resolved features. Across the predominant range of photon fluxes, the linewidth at $\Delta_{b}=0$ (blue line) is consistently narrower than that at $\Delta_{a}=0$ (red line), indicating that pumping at the $F=b$ resonance possesses an extra linewidth suppression capability, which fundamentally enables the magnetic resonance to be enhanced.
We explain this narrowing mechanism through the asymmetric absorption of $F=a$ and $F=b$ and its impact on the superoperator $\mathcal{G}_{\mathrm{WDA}}$ governing the relaxation of $|\tilde{\rho}_{1})$. Physically, for polarized atoms, the $F=a$ multiplet provides the dominant population imbalance that drives the coherence $|\tilde{\rho}_{1})$. Therefore, the overall relaxation of $|\tilde{\rho}_{1})$, which determines the resonance linewidth, is fundamentally dictated by the specific $\mathcal{G}_{\mathrm{WDA}}$ block corresponding to the $F=a$ level, i.e., $\mathcal{P}_{a}^{\{1\}}\mathcal{G}_{\mathrm{WDA}}\mathcal{P}_{a}^{\{1\}}$. Consistently, the optical pumping contribution $\mathcal{A}_{\mathrm{op}}]_{1,1}$ to this block originates from absorption of the $F=a$ level. This drives the conclusion that the direct light-induced relaxation incorporated into the linewidth is scaled by the weight factor $\mathcal{Q}_{a}$. When the pumping light is tuned near $\Delta_{b}=0$, $\mathcal{Q}_{a}$ attains values much smaller than 1 (see Fig. 2), manifesting as a practical suppression of decoherence during the optical pumping process. Conversely, the $F=a$ optical resonance exhibits $\mathcal{Q}_{a}>1$, resulting in a relative broadening of the linewidth. The $\mathcal{Q}_{a}$-scaling effect becomes particularly evident under intense light ($P\to 1$), where the pumping process dominates the linewidth behavior. In this case, the linewidth asymptotically approaches
$\Gamma_{2}=$ $\frac{1}{2I+1}A\Phi\mathcal{L}(\Delta_{a})+\frac{1}{2}\Gamma_{\mathrm{SD}},$ (63)
obtained by replacing the $F$-unresolved pumping rate in the HP counterpart [22] by $R_{\mathrm{op}}Q_{a}=A\Phi\mathcal{L}(\Delta_{a})$. This illustrates that the $F=b$ optical resonance essentially retards the growth of linewidth with increasing $\Phi$.
Sharing the same physical origin as the linewidth, the light modifies the Larmor frequency by the optical shift $\delta\omega_{\mathrm{L}}$ expressed as
$\delta\omega_{\mathrm{L}}$ $=\frac{\Delta_{a}}{\Gamma_{\mathrm{brd}}}\frac{A\Phi\mathcal{L}(\Delta_{a})}{2I+1}.$ (64)
While crossing zero at $\Delta_{a}=0$, detuning away from this point results in a residual shift. Fortunately, under favorable $\Gamma_{\mathrm{brd}}$ conditions, the optical shift can be attenuated by operating near the $F=b$ resonance where the perturbation to Larmor precession within the $F=a$ level is negligible. Notably, these QHP-specific theoretical predictions regarding the linewidth $\Gamma_{2}$ and light shift $\delta\omega_{\mathrm{L}}$ accord with previous experimental observations [5].
Furthermore, these features in physics are directly reflected in the observable $X$ and $Y$ signals, especially at large $\Phi$ as shown in Figs. 8(b) and 8(c). Although operating at $\Delta_{b}=0$ introduces a modest optical shift, it serves as a favorable trade-off for the overall magnetic resonance performance. When pumping at $\Delta_{b}=0$ (blue lines), the magnetic resonance spectra feature an enhanced signal amplitude compared to the $\Delta_{a}=0$ condition (red lines). We understand this enhancement behavior using the maximum resonance amplitude of the quadrature $Y$ as a representative metric. Due to its Lorentzian line shape, the resonance peak is evaluated by $Y_{\mathrm{max}}\sim P/\Gamma_{2}$, where $P$ acts as the fundamental source of $|\tilde{\rho}_{1})$. Coupling the linewidth suppression discussed above with the behavior of $P$ detailed in Sec. IV.2, we proceed to analyze $Y_{\mathrm{max}}$ across $\Phi$ domains. When $\Phi<\Phi^{\prime}_{\mathrm{eq}}$, the competing effects of $P$ and $\Gamma_{2}$ reach a close draw for $Y_{\mathrm{max}}$ between pumping at $\Delta_{a}=0$ and $\Delta_{b}=0$. In contrast, when $\Phi>\Phi^{\prime}_{\mathrm{eq}}$, the $F=b$ resonance not only maintains a higher $P$ to act as a powerful coherence source (enlarging the numerator), and meanwhile, its $\mathcal{Q}_{a}$-suppression effect protects $\Gamma_{2}$ from severe light broadening (lowering the denominator), thereby yielding a substantially enhanced magnetic resonance performance.
It should be noted that the physical demonstrations above primarily contrast the resonant pumping conditions where the $F$-resolved absorption and optical pumping behavior are most pronounced. In practical atomic magnetometer applications, while off-resonant pumping is still worth exploring for globally optimizing sensitivity, the $F=b$ near-resonance pumping configuration remains a preferred pathway due to the underlying physical mechanisms discussed above. In this section, the conclusions rooted in principles can be extended beyond the displayed $^{87}\text{Rb}$ system to other alkali-metal species.
V Conclusions
We developed a theoretical framework for optical pumping of alkali-metal atoms in buffer-gas cells operating in the QHP regime, where the collisional broadening $\Gamma_{\mathrm{brd}}$ is comparable to the GS hyperfine splitting $\omega^{\{\mathrm{g}\}}_{\mathrm{hf}}$. By incorporating the weighting factor $Q_{F}$ to account for asymmetric absorption, we derived the QHP-specific optical pumping superoperator within the Liouville-space master equation framework. Its $F$-resolved form established a compact relation linking to the HP counterpart, and facilitates both physical analysis and numerical evaluation in the QHP regime.
Using this unified approach, we demonstrated that QHP dynamics are governed by the combined effects of population redistribution across both $F$ multiplets and $m_{F}$ levels. We presented and discussed these $F$-pumping and $m_{F}$-pumping effects under different light polarizations. Furthermore, we examined the resulting steady-state populations that deviate from the standard STD and pointed out the validity of the STD approximation under these conditions.
Regarding observables with QHP corrections, we systematically evaluated three interconnected quantities under $F$-resolved pumping: the absorption cross section $\sigma$, the electron spin polarization $P$, and the intrinsic magnetic resonance linewidth $\Gamma_{2}$. The characteristic behaviors of absorption and polarization are highlighted by features such as the convergent integral area under linear polarization, alongside the distinct equal-absorption and equal-polarization points under circular polarization. Applying these insights to atomic magnetometry, we identified an optimal operating scheme, demonstrating that pumping near the $F=b$ resonance comprehensively enhances performance by suppressing the linewidth and amplifying the signal amplitude.
In summary, our study provides a profound understanding of the QHP-specific optical pumping and offers useful theoretical tools with practical applicability to atomic sensors (e.g. atomic magnetometers) operating under realistic buffer-gas pressures. The methodology developed herein also establishes a foundation for integrating further physical effects in the QHP regime, including atomic diffusion.
ACKNOWLEDGMENTS
This work is supported by XXXX.
DATA AVAILABILITY
The data that support the findings of this article are not publicly available. The data are available from the authors upon reasonable request.
Appendix A THE QHP CORRECTION IN LIOUVILLE SPACE
With the QHP correction Eq. (19), we offer detailed proof for the $F$-resolved form of superoperators $\mathcal{A}_{\mathrm{dp}}\mathcal{P}_{\mathrm{Z}}$ Eq. (22) and $\mathcal{A}^{\{\mathrm{e}g\}}_{\mathrm{p}}\mathcal{P}_{\mathrm{Z}}$ Eq. (27), and also show that this compact form are not available for hyperfine coherences between $F=a$ and $F=b$ multiplet.
To simplify the depopulation superoperator, we utilize the identities in Liouville space [9],
${X}^{\textcopyright}={X}^{\flat}-{X^{\dagger\sharp}},$ (65)
$(AB)^{\flat}=A^{\flat}B^{\flat},$ (66)
$(AB)^{\sharp}=B^{\sharp}A^{\sharp},$ (67)
with operators A,BA,\,B and $X$ in Hilbert space. Therefore, arising from the $F$-resolved $\delta H^{\{\mathrm{g}\}}$ in Eqs. (20) &\& (21), $\mathcal{A}_{\mathrm{dp}}$ Eq. (11) becomes
$\mathcal{A}_{\mathrm{dp}}=$ $\frac{i}{R_{\mathrm{op}}}\sum_{F}\left(\frac{\left(V^{\dagger}V\right)^{\flat}\Pi_{F}^{\flat}}{\Delta_{F}+i\Gamma_{\mathrm{brd}}}-\frac{\left(V^{\dagger}V\right)^{\sharp}\Pi_{F}^{\sharp}}{\Delta_{F}-i\Gamma_{\mathrm{brd}}}\right).$ (68)
The relation between projection operators in Hilbert and Liouville spaces is
$\Pi_{F}^{\flat}$ $=\mathcal{P}_{F}+\mathcal{P}_{F\to F^{\prime}},$ (69)
$\Pi_{F}^{\sharp}$ $=\mathcal{P}_{F}+\mathcal{P}_{F^{\prime}\to F}.$ (70)
Here, we define the projection operator of hyperfine coherence as ($F\neq F^{\prime}$)
$\mathcal{P}_{F\to F}$ $=\sum_{m_{F},m^{\prime}_{F}}\left|F,m_{F};F^{\prime},m^{\prime}_{F}\right)\left(F,m_{F};F^{\prime},m^{\prime}_{F}\right|,$ (71)
$\mathcal{P}_{\mathrm{hf}}$ $=\mathcal{P}_{a\to b}+\mathcal{P}_{b\to a}.$ (72)
These properties of projection operators enable us to derive
$\mathcal{A}_{\mathrm{dp}}\mathcal{P}_{\mathrm{Z}}=$ $\frac{i}{R_{\mathrm{op}}}\sum_{F}\left(\frac{\left(V^{\dagger}V\right)^{\flat}}{\Delta_{F}+i\Gamma_{\mathrm{brd}}}-\frac{\left(V^{\dagger}V\right)^{\sharp}}{\Delta_{F}-i\Gamma_{\mathrm{brd}}}\right)\mathcal{P}_{F},$ (73)
$\mathcal{A}_{\mathrm{dp}}\mathcal{P}_{\mathrm{hf}}=$ $\frac{i}{R_{\mathrm{op}}}\sum_{F}\left(\frac{\left(V^{\dagger}V\right)^{\flat}\mathcal{P}_{F\to F^{\prime}}}{\Delta_{F}+i\Gamma_{\mathrm{brd}}}-\frac{\left(V^{\dagger}V\right)^{\sharp}\mathcal{P}_{F^{\prime}\to F}}{\Delta_{F}-i\Gamma_{\mathrm{brd}}}\right).$ (74)
Here, the first line proves that, after the QHP correction, $\mathcal{A}_{\mathrm{dp}}\mathcal{P}_{\mathrm{Z}}$ acquires an $F$-resolved structure, allowing it to be decomposed into separate contributions from the $F=a$ and $F=b$ levels, in close analogy with $\delta H^{\{\mathrm{g}\}}$. In contrast, the second line indicates that the contributions from positive- and negative-frequency hyperfine coherences to the depopulation dynamics cannot be cast into a unified form in the same way. For further simplification, the weight factor $Q_{F}$ Eq. (23) is introduced
$Q_{F}\equiv\frac{\pi|\Omega_{\mathrm{I}}|^{2}\mathcal{L}(\Delta_{F})}{R_{\mathrm{op}}}=\frac{\mathcal{L}(\Delta_{F})}{x_{a}\mathcal{L}(\Delta_{a})+x_{b}\mathcal{L}(\Delta_{b})}.$ (75)
Therefore, we can employ the symmetry property of the dipole operator $\hat{\bm{e}}^{*}\cdot\bm{d}\bm{d}^{\dagger}\cdot\hat{\bm{e}}=1/2-K\mathbf{s}\cdot\mathbf{S}$ and obtain the compact form Eq. (22) expressed in terms of spin operators and projection operators.
We now turn to the optical excitation superoperator $\mathcal{A}^{\{\mathrm{eg}\}}_{\mathrm{p}}$. Following the same procedure as above, we focus on its $F$-resolved form Eq. (27). Starting from the definition Eq. (14), we obtain
$\mathcal{A}^{\{\mathrm{eg}\}}_{\mathrm{p}}\mathcal{P}_{\mathrm{Z}}=$ $\frac{i}{R_{\mathrm{op}}}\sum_{F}\left(\frac{V^{\dagger\sharp}V^{\flat}}{\Delta_{F}+i\Gamma_{\mathrm{brd}}}-\frac{V^{\dagger\sharp}V^{\flat}}{\Delta_{F}-i\Gamma_{\mathrm{brd}}}\right)\mathcal{P}_{F}.$ (76)
As in the similar case of Eq. (74), this form also does not extend to the hyperfine coherence. Here, we utilize the property ${B}^{\mathrm{T}}\otimes A=A^{\flat}B^{\sharp}$ [9]. With the factor $Q_{F}$, $A^{\{\mathrm{eg}\}}_{\mathrm{p}}\mathcal{P}_{\mathrm{Z}}$ is expressed as Eq. (27). In the HP limit $Q_{F}\to 1$, the excitation dynamics becomes $F$-unresolved, and the superoperator reduces to
[…]
In summary, the QHP corrections derived for the various superoperators consistently follow the same $Q_{F}$-weighted and $F$-resolved formalism, thereby establishing a clear connection to the traditional HP limit. This structural equivalence is essential for constructing a compact expression of the optical pumping superoperator $\mathcal{A}_{\mathrm{op}}$.
Appendix B FREQUENCY BLOCKS IN ZEEMAN SUBSPACE
In Liouville-space calculations for the master equation, an appropriate ordering of the basis states is essential for exposing the block structure of large superoperators. Such a structure not only improves computational efficiency but also highlights the underlying physical organization of the dynamics. A block group scheme is constructed in Zeeman subspace, i.e., $\mathrm{span}\{|Fm_{F},Fm^{\prime}_{F})\}$ according to the characteristic evolution frequencies of the density-matrix elements [22].
The projection operators of the GS $k$th-order Zeeman coherence are introduced as
$\mathcal{P}^{\{k\}}_{\mathrm{Z}}$ $=\mathcal{P}^{\{k\}}_{a}+\mathcal{P}^{\{k\}}_{b}$, (78)
$\mathcal{P}^{\{k\}}_{a}$ $=\sum_{m_{a}}\left|a,{m}_{a}+k;a,{m}_{a}\right)\left(a,{m}_{a}+k;a,{m}_{a}\right|,$ (79)
$\mathcal{P}^{\{k\}}_{b}$ $=\sum_{m_{b}}\left|b,{m}_{b}-k;b,{m}_{b}\right)\left(b,{m}_{b}-k;b,{m}_{b}\right|.$ (80)
where $\left|F,m_{F};F^{\prime},m^{\prime}_{F^{\prime}}\right)$ is the basis of $|{\rho})$ in Liouville space,
$|F,m_{F};F^{\prime},m^{\prime}_{F})=|F,m_{F}\rangle\langle F^{\prime},m^{\prime}_{F}|.$ (81)
[…]
$\mathcal{X}$ $=\sum_{m,n}[\mathcal{X}]_{m,n},$ (82)
$[\mathcal{X}]_{m,n}$ $=\mathcal{P}^{\{m\}}_{\mathrm{Z}}[\mathcal{X}]\mathcal{P}^{\{n\}}_{\mathrm{Z}}$, (83)
where $\mathcal{X}]_{m,n}$ physically couples $n$th-order to $m$th-order Zeeman coherence subspace.
To simplify the full dynamical evolution of $\left|{\rho}\right)$, we neglect the coherences between the hyperfine multiplets and focus on mechanisms in the Zeeman Subspace. This is because in the absence of microwave driving, the hyperfine coherences will decay within a timescale typically on the order of approximately milliseconds or even shorter [6].